Macrohard Corporation is a new business based out of Puerto Rico that is considering a plan to manufacture and market cellphones in Peurto Rico starting in 2022. However, before it decides to take the plunge, it needs to understand the feasibility of breaking even. Macrohard is planning to sell the cellphones to wholesalers at $25. The electronic circuitry costs $7.50, CORNING supplies the screen for $3.75, and the battery costs $1.25. Additionally, the cost of a fabrication facility is $8.5 million, and Macrohard estimates that it will need to spend $2 million on sales and marketing expenses in order to achieve its objectives, and $2 million on general and administrative expenses. Calculate the break-even volume, and explain how feasible it is for Macrohard to break even.  
Answer #1

 Answer:-

Total fixed cost= FC= 2,000,000+2,000,000+8,500,000= 12,500,000

Total variable cost= VC= 7.5+3.75+1.25= 12.5

Selling price= SP= 25

Break even point= FC/(SP-VC)= 12,500,000/(25-12.5)= 1,000,000

Hence 1 million cellphones needs to be sold to break even. Any volume below this and the company wont turn a profit.


answered by: Sehaj Kaur